# Arbitrarily Applicable Same/Opposite Relational Responding with NARS

### v1

@@ 2.2 Same/Opposite Relational Responding
Within RFT, relational responding includes diverse relational frames such as opposite, distinction, comparison, and hierarchy [2]. The relational frame of opposition involves responding to stimuli as functionally opposite based on contextual cues. Combining relational frames (e.g., SAME and OPPOSITE) through mutual and combinatorial entailments enables sophisticated cognitive behaviors like categorization, discrimination, and generalized symbolic reasoning.
@@ 3 Implementation in NARS
3.1 Matching-to-Sample Task in Narsese
The Matching-to-Sample (MTS) procedure was explicitly represented using Narsese temporal statements, capturing relationships among stimuli, locations, and relational context. A representative trial is encoded as follows:
@@ 4 Experimental Setup
Figure 1: The Matching-to-Sample task used in the present study. Pre-training phases are excluded from this figure. The figure illustrates Phases 2–3, with underlined options indicating correct responses.
@@ 6 Discussion
However, this study explicitly has limitations. The relational structures explored were relatively simple and abstract. Future work should explicitly examine more complex relational networks and multiple relational frames, potentially integrating real-world sensorimotor data to enhance ecological validity and applicability in naturalistic contexts. Additionally, addressing scalability and generalizability explicitly remains critical for practical AGI applications, where relational reasoning must operate robustly amidst noisy, ambiguous, or incomplete real-world data.
In conclusion, the present findings explicitly represent a meaningful advancement in computational cognitive modeling, illustrating concretely how psychological theories such as RFT can explicitly inform and enrich AGI architectures. This integration explicitly advances AGI methodologies and clearly illustrates how arbitrary relational contexts underpin human symbolic reasoning, contributing meaningfully toward more robust, flexible, and human-like artificial intelligence.

### v2

@@ 2.1 Arbitrarily Applicable Relational Responding and Relational Frame Theory
2.2 Matching-to-sample task in RFT
The Matching-to-Sample (MTS) task is a widely-used experimental paradigm in the study of relational responding, particularly within the framework of Relational Frame Theory (RFT). In a typical MTS procedure, participants are first presented with a sample stimulus and then choose from comparison stimuli based on relational rules explicitly taught or inferred from context. A relational contextual cue (such as SAME or OPPOSITE) can also be presented above the sample stimulus, specifying the relational response required. Successful performance relies on participants’ ability to derive relational responses, illustrating mutual entailment and combinatorial entailment. The MTS paradigm thus effectively assesses participants’ capacity for arbitrarily applicable relational responding, especially regarding stimulus equivalence and relational generalization. Figure 1 illustrates a typical Matching-to-Sample task scenario used in RFT research.
Figure 1: The Matching-to-Sample task used in the present study. Pre-training phases are excluded from this figure. The figure illustrates experimental phases 2–3, with underlined options indicating correct responses.
@@ 2.3 Same/Opposite Relational Responding
Within RFT, relational responding includes diverse relational frames such as sameness, opposite, distinction, comparison, and hierarchy [2]. The relational frame of opposition involves responding to stimuli as functionally opposite based on contextual cues. Combining relational frames (e.g., SAME and OPPOSITE) through mutual and combinatorial entailments enables sophisticated cognitive behaviors like categorization, discrimination, and generalized symbolic reasoning.
@@ 3 Implementation in NARS
3.1 Same/Opposite Matching-to-Sample Task in Narsese
The Matching-to-Sample (MTS) procedure was explicitly represented using Narsese temporal statements, capturing relationships among stimuli, locations, and relational context. This type of encoding has been described extensively elsewhere [3, 4]. A representative trial is encoded as follows:
@@ 6 Discussion
However, this study has limitations. The relational structures explored were relatively simple and abstract. Future work should explicitly examine more complex relational networks and multiple relational frames, potentially integrating real-world sensorimotor data to enhance ecological validity and applicability in naturalistic contexts. Additionally, addressing scalability and generalizability explicitly remains critical for practical AGI applications, where relational reasoning must operate robustly amidst noisy, ambiguous, or incomplete real-world data.
A further limitation is the reliance on empirical validation rather than formal proofs of completeness or consistency of the acquired-relations mechanism. While the findings empirically demonstrate robust relational responding, formal analyses of theoretical expressiveness or soundness were beyond the study’s scope. Future work could explicitly investigate these formal properties to enhance theoretical rigor.
In conclusion, the present findings explicitly represent a meaningful advancement in computational cognitive modeling, illustrating concretely how psychological theories such as RFT can explicitly inform and enrich AGI architectures. This integration explicitly advances AGI methodologies and clearly illustrates how arbitrary relational contexts underpin human symbolic reasoning, contributing meaningfully toward more robust, flexible, and human-like AI.